OpenStudy (anonymous):

If you draw one card from a standard deck, what is the probability that you draw a king or queen?

OpenStudy (bahrom7893):

There are 4 kings and 4 queens.

OpenStudy (bahrom7893):

There are 52 cards in a deck..

OpenStudy (bahrom7893):

So: p=8/52=2/13